# Intergroup violence in bursts or fizzles

### v9

@@ 8 Composition of indicators
Turmoil. We subsumed the following behaviors of members of the opponent group under the heading of turmoil for the focal group. First, we coded approaching when opponents moved closer to focal group members. We also indicated whether actors moved energetically (jumping/skipping) and whether they ran. Second, aggressive behaviors when actors in the opponents’ group moved body parts in a way that signaled provocation, hostility and/or a readiness to attack. We used the following modifiers (sub-codes) to specify aggressing behavior: fighting gestures; pulling off clothing (jackets or vests); pulling up pants; pointing at opponents; provocative gesturing with fingers or hands (as an invitation to engage); bending forward (head and/or chest bending toward opponent); encroaching (invading opponents’ personal space through using or damaging objects belonging to them); and teasing (invading opponents’ body space in a teasing/ridiculing way that does not inflict bodily harm, such as lightly hitting, stroking, or patting the opponent’s body). Third, involuntary behavior that increases the vulnerability of opponents, because it tends to agitate (signaling opportunity) and provoke violence. This concerned stumbling and falling to the ground. Finally, violence committed by members of the opponent group. Because we do not know the psychological impact of the various kinds of turmoil, we could not construct a ratio level scale. Instead, we calculated the total level of turmoil from the beginning of the clip until a focal group’s maximum participation in violence by noting the duration of each instance of turmoil, and multiplying it by the number of opponents involved, i.e., the surface area under the turmoil curve in Fig. S10, S13, and S14.
@@ 9 Examples from videos
Graphs of behavior on the timeline enable to analyze and compare the coded videos. Each graph shows the intervals and their durations of (collective) violence, turmoil, synchrony and deescalation (see Fig. S10, S13, S14).
To illustrate, we discuss how coded video data was turned into a graph for clip 26. Fig. S7-9 feature screenshots of the clip, and Fig. S10 displays the graph. This clip features a confrontation of a dyad (focal group) with a triad (opponents) in a restaurant. Violence started at approximately 14 seconds into the video and quickly became collective until approximately 28 seconds. Four instances of turmoil preceded the violence, starting with one of the members of the opponent group gesturing aggressively and walking around in a jumpy, energetic way. The members of the focal dyad walked synchronously toward and away from the opponent in three shorter and one relatively longer stretch of time. Fig. S7 captures a moment of synchrony; the members of the dyad were making fighting gestures and stood aligned, each stretching out one leg with their feet pointing toward the opponent, and the other leg standing backward. Fig. S8 shows the moment when the dyad’s level of agitation reached a critical level; when one of the focal actors dodged backward, avoiding being hit, his companion was about to deliver a blow to the head of the opponent, which made the latter fall down. Fig. S9 displays the second member of the opponents’ triad (soon joined by the third) rushing into the restaurant toward the focal dyad, near his fallen companion. At this point, both groups were engaged in collective violence. Fig. S8 also shows bystanders (who can be seen sitting at a table in Fig. S7) and employees standing near the counter. At 28 seconds, an employee took deescalatory action for the first time. At that point, the collective violence had stopped already and erupted again for a shorter moment approximately 2 seconds later. The clip ends when two triad members carry their fallen group member outside. By then, the focal dyad had already left the scene.
@@
Figure 10: Graph of clip 26 with a burst of collective violence in a dyad.
Fig. S11-S13 provide another illustration. Fig. S13 is the graph of clip 86; it shows three outbreaks of collective violence in a group of five members who faced a single opponent. In the first instance of collective violence (not a burst by our 2 seconds criterion), starting at approximately 60 seconds into the clip and lasting for approximately 30 seconds, four members participated. Three of them immediately followed each other, and the fourth joined after approximately 2 seconds. Prior to the outbreak of violence, the opponent agitated focal group members by walking back and forth and by posturing aggressively. Fig. S11 captures the trigger moment of agitation; i.e., while one member of the focal group was engaged in aggressive posturing and a group member tried to deescalate, the opponent was about to strike another focal group member. Fig. S12 shows the situation 4 seconds later, when three members of the focal group were attacking.
@@ 11 Overview of cases and descriptive statistics
Table SLABEL:table:long: Overview of all groups and their sizes; occurrence of collective violence (c.v.); maximum participation in violence ($N_{C}$); time lag between the first participant and maximum participation in collective violence; burst; levels of turmoil ($T$) and of synchrony at first violence outbreak; number of unconditional defectors ($N_{D}$). In group 8_queensday, two focal group members disappeared from view already before the first moment of violence had started; out of the remaining 12, 6 participated in collective violence.
@@ 6 Burst 23 (39) 0 1
Table 2: Descriptive statistics of indicators used in the analysis ($\mathcal{N}$=59). Note: Levels of synchrony prior to maximum violence includes an outlier of 303.42. Without outlier, mean = 11.29; s.d. = 18.96; max = 78.75. We excluded this outlier in the calculations reported.
@@ 12.1 Mean field analysis
with $\hat{S_{i}}$ and $\hat{S_{j}}\in\{-\Delta,\Delta\}$, and taking into account the row-normalization of the adjacency matrix ($\sum_{j}^{n}w_{ij}=1$) [10].
@@
This result without unconditional defectors was inferred in an earlier study [10]. Here, we also deal with unconditional defectors in proportion $p>0$. Accordingly, we define $M_{cc}$ as the average spin value of the conditional cooperators and $M_{ud}$ as the average spin value of the unconditional defectors. Note that $M_{ud}=S^{-}$. We assume that the unconditional defectors are homogeneously distributed across the network. Accordingly, the mean field equation becomes
@@ 12.3 Solving the Ising model computationally
In simulations, $p_{c}$ is less precise in smaller networks (Table S3) and the prediction for the triad is empirically false, whereas the mean field is correct: one unconditional defector does not prevent cooperation of the other two. Also one empirical group of six with two unconditional defectors had a burst, which did not happen in the simulations.
@@
JB made the asymmetric Ising model, and wrote the software and the paper. DW collected, interpreted, and analyzed the data. BM did the mean field analysis. JB and DW wrote the Supplementary Material.

### v10

@@ 9 Composition of indicators
Turmoil. We subsumed the following behaviors of members of the opponent group under the heading of turmoil for the focal group. First, we coded approaching when opponents moved closer to focal group members. We also indicated whether actors moved energetically (jumping/skipping) and whether they ran. Second, aggressive behaviors when actors in the opponents’ group moved body parts in a way that signaled provocation, hostility and/or a readiness to attack. We used the following modifiers (sub-codes) to specify aggressing behavior: fighting gestures; pulling off clothing (jackets or vests); pulling up pants; pointing at opponents; provocative gesturing with fingers or hands (as an invitation to engage); bending forward (head and/or chest bending toward opponent); encroaching (invading opponents’ personal space through using or damaging objects belonging to them); and teasing (invading opponents’ body space in a teasing/ridiculing way that does not inflict bodily harm, such as lightly hitting, stroking, or patting the opponent’s body). Third, involuntary behavior that increases the vulnerability of opponents, because it tends to agitate (signaling opportunity) and provoke violence. This concerned stumbling and falling to the ground. Finally, violence committed by members of the opponent group. Because we do not know the psychological effect sizes of the various kinds of turmoil, we could not construct a ratio level scale. Instead, we calculated the total level of turmoil from the beginning of the clip until a focal group’s maximum participation in violence by noting the duration of each instance of turmoil, and multiplying it by the number of opponents involved, i.e., the surface area under the turmoil curve in Fig. S9, S12, and S13.
@@ 10 Examples from videos
Graphs of behavior on the timeline enable to analyze and compare the coded videos. Each graph shows the intervals and their durations of (collective) violence, turmoil, synchrony and deescalation (see Fig. S9, S12, S13).
To illustrate, we discuss how coded video data was turned into a graph for clip 26. Fig. S6-8 feature screenshots of the clip, and Fig. S9 displays the graph. This clip features a confrontation of a dyad (focal group) with a triad (opponents) in a restaurant. Violence started at approximately 14 seconds into the video and quickly became collective until approximately 28 seconds. Four instances of turmoil preceded the violence, starting with one of the members of the opponent group gesturing aggressively and walking around in a jumpy, energetic way. The members of the focal dyad walked synchronously toward and away from the opponent in three shorter and one relatively longer stretch of time. Fig. S6 captures a moment of synchrony; the members of the dyad were making fighting gestures and stood aligned, each stretching out one leg with their feet pointing toward the opponent, and the other leg standing backward. Fig. S7 shows the moment when the dyad’s level of agitation reached a critical level; when one of the focal actors dodged backward, avoiding being hit, his companion was about to deliver a blow to the head of the opponent, which made the latter fall down. Fig. S8 displays the second member of the opponents’ triad (soon joined by the third) rushing into the restaurant toward the focal dyad, near his fallen companion. At this point, both groups were engaged in collective violence. Fig. S7 also shows bystanders (who can be seen sitting at a table in Fig. S6) and employees standing near the counter. At 28 seconds, an employee took deescalatory action for the first time. At that point, the collective violence had stopped already and erupted again for a shorter moment approximately 2 seconds later. The clip ends when two triad members carry their fallen group member outside. By then, the focal dyad had already left the scene.
@@
Figure 9: Graph of clip 26 with a burst of collective violence by a dyad.
Fig. S10-S12 provide another illustration. Fig. S12 is the graph of clip 86; it shows three outbreaks of collective violence in a group of five members who faced a single opponent. In the first instance of collective violence (not a burst by our 2 seconds criterion), starting at approximately 60 seconds into the clip and lasting for approximately 30 seconds, four members participated. Three of them immediately followed each other, and the fourth joined after approximately 2 seconds. Prior to the outbreak of violence, the opponent agitated focal group members by walking back and forth and by posturing aggressively. Fig. S10 captures the trigger moment of agitation; i.e., while one member of the focal group was engaged in aggressive posturing and a group member tried to deescalate, the opponent was about to strike another focal group member. Fig. S11 shows the situation 4 seconds later, when three members of the focal group were attacking.
@@ 12 Overview of cases and descriptive statistics
Table SLABEL:table:long: Overview of the groups. Occurrence of collective violence (c.v.); maximum participation in violence ($N_{C}$); time lag between the first participant and maximum participation in collective violence; burst; levels of opponents-produced turmoil ($T$) and of synchrony at first violence outbreak; number of unconditional defectors ($N_{D}$). In group 8_queensday, two focal group members disappeared from view already before the first moment of violence had started; out of the remaining 12, 6 participated in collective violence.
@@ 6 Burst 23 (39) 0 1
Table 2: Descriptive statistics. Note: Levels of synchrony prior to maximum violence includes an outlier of 303.42. Without outlier, mean = 11.29; s.d. = 18.96; max = 78.75. We excluded this outlier in the calculations reported.
@@ 13.1 Mean field analysis
with $\hat{S_{i}}$ and $\hat{S_{j}}\in\{-\Delta,\Delta\}$, and taking into account the row-normalization of the adjacency matrix ($\sum_{j}^{n}w_{ij}=1$).
@@
This result without unconditional defectors was inferred in an earlier study (author(s)). Here, we also deal with unconditional defectors in proportion $p>0$. Accordingly, we define $M_{cc}$ as the average spin value of the conditional cooperators and $M_{ud}$ as the average spin value of the unconditional defectors. Note that $M_{ud}=S^{-}$. We assume that the unconditional defectors are homogeneously distributed across the network. Accordingly, the mean field equation becomes
@@ 13.3 Solving the Ising model computationally
In simulations, $p_{c}$ is less precise in smaller networks (Table S3) and the prediction for the triad is empirically false, whereas the mean field is correct: one unconditional defector does not prevent cooperation of the other two. Also one empirical group of six with two unconditional defectors had a burst, which did not happen in the simulations (Fig 15).
@@
Figure 15: Color online. For simulated networks with sizes $3\leq n\leq 100$, the minimum number of unconditional defectors who prevent a burst is plotted. The red dot (left, closest to the bottom) marks the triad (1 defector prevents a burst of 2), which conflicts with the empirical data (1 defector and a burst of 2).
@@
JB made the asymmetric Ising model, and wrote the software and the paper. DW collected, interpreted, and analyzed the data. BM did the mean field analysis. JB and DW wrote the Supplementary Material. No AI has been used in this research project.